Let f(x) = 3x + 2 and g(x) = 7x + 6 of f*g.
Then describe the domain of f*g (isn't it all real numbers)
And what about this one:

Let f(x) = 3x - 6 and g(x) = x - 2. Find f/g.
Then describe the domain of f/g.
Thanks in advance!

Answer by ewatrrr(24785)

 
Hi,
f(x) = 3x + 2 and g(x) = 7x + 6
f*g = (3x+2)(7x+6) = 21x^2 + 32x + 12 |Yes, domain is all real numbers
f(x) = 3x - 6 and g(x) = x - 2
f/g = (3x-6)/(x-2) |Domain is all real numbers, except x = 2
CANNOT divide by zero